# From Simple to Complex: Knowledge Transfer in Safe and Efficient Reinforcement Learning for Autonomous Driving

### v1

@@ Abstract
A safe and efficient decision-making system is crucial for autonomous vehicles. However, the complexity of driving environments limit the effectiveness of many rule-based and machine learning-based decision-making approaches. The introduction of Reinforcement Learning in autonomous driving presents a promising solution to these challenges, although concerns about safety and efficiency during training remain major obstacles to its widespread application. To address these concerns, we propose a novel framework named Simple to Complex Collaborative Decision. First, we rapidly train the teacher model using the Proximal Policy Optimization algorithm in a lightweight autonomous driving simulation environment. In the more complex simulation environment, the teacher model intervenes when the student agent exhibits sub-optimal behavior by assessing the value of actions to avert dangerous situations. Next, we developed an innovative algorithm called Adaptive Clipping Proximal Policy Optimization. It trains using a combination of samples generated by both the teacher and student policies and applies dynamic clipping strategies based on sample importance, enabling the algorithm to utilize samples from diverse sources more efficiently. Additionally, we employ the KL divergence between the teacher’s and student’s policies as a constraint for policy optimization to facilitate the student agent’s rapid learning of the teacher’s policy. Finally, by adopting an appropriate weaning strategy to gradually reduce teacher intervention, we ensure that the student agent can fully explore the environment independently during the later stages of training. Simulation experiments in highway lane-change scenarios demonstrate that, compared to baseline algorithms, our proposed framework not only improves learning efficiency and reduces training costs but also significantly enhances safety during training.
@@ 1 Introduction
Due to the complexity and unpredictability of driving environments, approximately 94% of traffic accidents are caused by suboptimal decisions made by human drivers (Alvaro et al., 2018). Suboptimal driving decisions not only threaten traffic safety but also reduce overall traffic efficiency (Lee et al., 2004). The rapid advancement of autonomous driving technology has attracted significant attention due to its potential to improve both traffic safety and efficiency. Autonomous driving systems generally consist of a perception layer, a decision-making layer, and a planning and control layer, with the decision-making layer commonly referred to as the "brain" of the system (Chen et al., 2017). Therefore, ensuring the safety and accuracy of decision-making algorithms is crucial for the overall performance of autonomous driving systems (Li et al., 2017). Currently, decision-making methods for autonomous vehicles are broadly classified into two categories: knowledge-driven methods and data-driven methods. Knowledge-driven methods, such as Hierarchical State Machines (HSM) (Patz et al., 2008), Expert Systems (ES) (Ferguson et al., 2008), and Finite State Machines (FSM) (Buehler et al., 2009), are characterized by their rigorous logic and high interpretability (Bae et al., 2020). However, these methods are constrained by their strong reliance on existing knowledge, which limits their ability to address scenarios beyond their knowledge base (Bianco-Vega et al., 2005; Mozina et al., 2008). With the rise of Deep Learning, data-driven methods like End-to-End (E2E) Learning (Bojarski et al., 2016), Imitation Learning (IL) (Hawke et al., 2020), and RL (Sallab et al., 2017; Zhang et al., 2022b) have become prominent research topics in autonomous driving. Data-driven methods offer robust self-learning capabilities and environmental adaptability, making them well-suited for addressing decision-making needs in various scenarios, particularly through the use of RL algorithms. RL algorithms optimize driving strategies through interactions with the environment, enabling better adaptation to complex driving scenarios (Zhu and Zhao, 2021). Despite the impressive performance of RL in fields such as game artificial intelligence (AI) (Lample and Chaplot, 2017; Ye et al., 2020), financial transactions (Deng et al., 2016; Pendharkar and Cusatis, 2018), and energy management (Yu et al., 2019; Kuznetsova et al., 2013), significant challenges remain. Safety risks and training inefficiencies hinder its widespread application (Liu et al., 2022), particularly in safety-critical areas like autonomous driving and robotic control.
@@
RL, similar to human behavior, can autonomously explore environments and optimize strategies based on environmental feedback. However, traditional RL algorithms inevitably encounter dangers before learning how to avoid them (Chow et al., 2018), and occasionally they must engage in risky actions to learn effective strategies. The cost of allowing a vehicle to undertake dangerous actions in real-world scenarios is prohibitively high, making it challenging to apply traditional RL algorithms in autonomous driving. In human learning, especially when dealing with dangerous situations, we do not rely solely on trial and error to acquire knowledge (Peng et al., 2022b). Instead, we seek guidance from a teacher to ensure safety and efficiency during the learning process. The teacher not only demonstrates correct actions when necessary but also directly corrects the student’s mistakes to avert dangerous situations. For example, when a student learns to drive, an experienced instructor accompanies the student and can take over control of the vehicle if the student makes a poor decision due to inexperience, thereby preventing an accident. Consequently, numerous researchers have incorporated the TSF (Zimmer et al., 2014; Liu et al., 2023) into RL, enabling student agents to learn how to manage complex scenarios safely and efficiently through both autonomous exploration and teacher demonstrations. In the TSF, a high-performance neural network controller or a human expert is integrated as the teacher policy into the student agent’s learning process. At each step, the teacher intervenes in the student’s actions when specific intervention criteria are met. When an intervention occurs, the teacher’s actions replace the student’s actions in the replay buffer as training data, which is subsequently used to optimize the student’s model via RL algorithms. The human-in-the-loop TSF requires continuous monitoring and intervention by experts throughout the training process, resulting in substantial labor costs. Moreover, the teaching effectiveness of human experts is easily affected by subjective factors such as emotions and mental states. To address these issues, most TSF approaches first train a high-performance teacher model using neural networks, enabling the student agent to learn as many correct behaviors as possible from the teacher’s demonstrations. However, in many application domains, such as robotic control and autonomous driving, acquiring a high-performance teacher model in advance can be prohibitively expensive. Furthermore, in current TSF methods, if the teacher fails to provide high-quality demonstrations, the student agent may learn incorrect strategies, ultimately limiting its performance due to the teacher’s flawed approach (Xue et al., 2023). At the same time, regardless of whether the TSF ultimately selects the teacher’s or the student’s action at each step, the RL algorithm can only utilize the Markov Decision Process (MDP) corresponding to the chosen action as the training sample. The inability to exploit another action sample leads to considerable data wastage.
@@
In summary, traditional RL algorithms not only involve high training costs but also fail to guarantee the safety of agents during training. Although standard TSF methods enable the student agent to learn safely and efficiently, they incur substantial costs in acquiring a high-performance teacher, and the student’s final performance is limited by the teacher’s capabilities. Furthermore, these methods cannot optimize the policy by simultaneously utilizing samples generated by both the teacher and student models. Therefore, in this paper, we propose a novel TSF called the S2CD framework, which effectively facilitates knowledge transfer from teacher to student. Unlike other TSFs, the S2CD framework first trains the teacher model in a lightweight simulation environment with similar scenarios, thereby simplifying and accelerating the training process. Subsequently, the teacher guides the student to learn safely and efficiently in more complex simulation or real-world environments. We also theoretically demonstrated that teacher guidance can effectively enhance both the performance ceiling and safety ceiling of the student during training. Additionally, we improved upon the PPO algorithm used in S2CD and proposed the ACPPO algorithm. This algorithm simultaneously utilizes data from both the teacher and student models to significantly improve sample efficiency. It also adaptively adjusts the clipping factor based on sample importance, enhancing learning efficiency. During the policy optimization process, we update the neural network by constraining the KL divergence between the teacher’s and student’s policies, which facilitates the rapid convergence of the student policy to the teacher policy. Finally, S2CD implements an appropriate weaning strategy to ensure that the student agent can independently explore the environment in the later stages, thereby overcoming the limitations of the teacher’s capabilities. The main contributions of this paper are summarized as follows:
@@
  • This paper presents a novel TSF called S2CD, which rapidly trains a teacher model in a lightweight simulation environment and subsequently uses that teacher to guide the student agent in learning safely and efficiently in more complex simulated or real-world environments.
@@
  • This paper proposes the ACPPO algorithm, which is integrated within the S2CD framework. The ACPPO algorithm simultaneously leverages data from both the teacher and student models and adaptively adjusts the clipping factor based on the importance of each sample, thereby improving sample efficiency.
@@
  • This paper introduces the use of the KL divergence between the teacher’s and student’s policies as a constraint in policy updates, transforming it into an unconstrained problem using the Lagrangian method. This approach enables the student’s policy to rapidly approximate the teacher’s policy, thereby enhancing learning efficiency.
@@
  • This paper examines the teacher’s intervention and weaning strategies within the S2CD framework, where appropriate intervention moments are selected based on action values. In the later stages, the probability of teacher demonstrations and interventions is gradually decreased to ensure the student model’s capacity for independent exploration, thereby overcoming limitations due to the teacher’s suboptimal performance.
@@ 2.1 Knowledge Transfer
While RL has been widely applied across various fields and has demonstrated impressive results, most research has primarily focused on simulation experiments. However, in the context of autonomous driving, the ultimate goal is real-world application, where safety and efficiency during training process are critical factors in assessing algorithm performance. Training standard RL algorithms on real vehicles would not only risk substantial damage but also incur significant time costs. Consequently, effectively transferring knowledge from simulation-based training to agents operating in real-world or more complex simulation environments has become a significant focus of research. For example, Qiao et al. (2018) and Song et al. (2021) employed Curriculum Learning (CL) strategies to achieve notable performance in urban intersection and overtaking tasks, respectively. CL enhances model learning effectiveness by gradually increasing task difficulty, instead of requiring models to tackle complex tasks directly (Bengio et al., 2009). However, CL requires meticulously designed multi-level tasks. Furthermore, the progression from simpler to more complex tasks can increase training time and computational resource demands. Transfer Learning (TL) is another knowledge transfer technique that effectively reduces the training costs of the target task by applying knowledge acquired from the source task to the target task. For instance, Shu et al. (2021) combined RL with TL to transfer decision-making knowledge from one driving task to another in an intersection environment. Similarly, Akhauri et al. (2020) utilized simulated accident data for TL, enabling the model to generalize more rapidly and effectively to real-world scenarios. However, if the source and target tasks are not sufficiently related, the performance of TL may deteriorate, potentially resulting in "negative transfer" (Zhang et al., 2022a). Another approach to knowledge transfer is the use of Digital Twins (DT). By creating an accurate digital replica of a physical entity or system, we can map data from the physical world to the virtual world (Niaz et al., 2021), enabling real-time monitoring, simulation, analysis, and optimization of physical entities. Voogd et al. (2023) employed DT technology to combine virtual and real-world data on vehicle dynamics and traffic scenarios, effectively bridging the gap between simulation and reality. Wang et al. (2024) designed and implemented an E2E DT system for autonomous driving. This system captures real-world traffic information, processes it in the cloud to create a digital twin model, and then utilizes this model to provide route planning services for autonomous vehicles. Despite DT technology effectively bridging the gap between simulation training and real-world applications, accurate simulation processes remain indispensable. Designing simulation programs that accurately mirror or fully replicate real-world conditions is both challenging and costly (Yun and Park, 2021). Moreover, pre-training a high-performance model in a high-fidelity simulation environment is resource-intensive.
@@ 2.2 Safe RL
RL methods can enhance an agent’s ability to avoid dangers by learning from previous training failures and developing efficient behavioral strategies to adapt to complex environments. However, the application of these methods in safety-critical domains continues to present substantial safety challenges (Garcıa and Fernández, 2015). Ensuring policy safety throughout the training process and preventing agents from encountering dangerous situations are key to expanding the applicability of RL to more fields. Consequently, many researchers are focusing on incorporating safety constraints into RL algorithms to mitigate unsafe behaviors during both the learning and execution phases. For instance, methods like PPO-Lag (Stooke et al., 2020) and SAC-Lag (Ha et al., 2021) utilize the Lagrangian method, which simultaneously updates both the policy and the Lagrange multiplier through primal-dual optimization, thereby transforming the safety-constrained optimization problem into an unconstrained one. Peng et al. (2022a) proposed the Separated Proportional Integral Lagrangian (SPIL) algorithm, which mitigates oscillations and reduces conservatism in RL policies during car-following scenarios. Zhang et al. (2024) introduced cognitive uncertainty as a constraint in Lagrangian Safe RL algorithms, aiming to enhance both exploration and safety performance in autonomous vehicles. While the Lagrangian method is effective in handling safety constraints, it lacks a clear mechanism to prevent hazardous events (Li et al., 2022b). Thus, agents may still engage in unsafe exploration during training. Consequently, many researchers have applied rule-based constraints (Cao et al., 2022; Likmeta et al., 2020) or risk estimation methods (Mo et al., 2021; Li et al., 2022a) to Safe RL in autonomous driving, aiming to ensure a minimum level of policy safety performance. Although these methods can prevent a significant number of unsafe actions, their effectiveness heavily depends on the quality of the rule design or risk assessment models. To further reduce unsafe events, researchers have shifted their focus toward Safe RL approaches based on the TSF. Within the TSF, teacher policies can be classified into human expert strategies and well-trained neural network strategies. For example, Wu et al. (2023) and Huang et al. (2024) employed human experts to mentor AI agents. During the student agent’s interaction with the environment, the human expert intervenes in dangerous situations and demonstrates the correct actions, thereby ensuring driving safety throughout the student’s training process. While Safe RL methods involving human expert intervention significantly improve the safety of the exploration process, they incur high labor costs. Moreover, the teaching performance of human experts is highly influenced by subjective factors, leading to instability in the final model performance. To reduce labor costs, many researchers have employed well-trained neural network models as "teachers" to guide student agents during the learning process. For instance, Peng et al. (2022b) and Xue et al. (2023) employed RL algorithms to obtain a well-trained teacher model, when the conditions defined by a switching function, which determines whether to intervene, are met, the teacher takes over the student’s actions to demonstrate correct operations and avoid dangerous situations. In our previous work (Zhou et al., 2024), we employed the SOAR cognitive architecture as a teacher to guide the student agent toward efficient and safe learning, while also offering a certain degree of interpretability.

### v3

# Knowledge Transfer from Simple to Complex: A Safe and Efficient Reinforcement Learning Framework for Autonomous Driving Decision-Making

@@ Abstract
A safe and efficient decision-making system is crucial for autonomous vehicles. However, the complexity of driving environments limits the effectiveness of many rule-based and machine learning approaches. Reinforcement Learning, with its robust self-learning capabilities and environmental adaptability, offers a promising solution to these challenges. Nevertheless, safety and efficiency concerns during training hinder its widespread application. To address these concerns, we propose a novel RL framework, Simple to Complex Collaborative Decision (S2CD). First, we rapidly train the teacher model in a lightweight simulation environment. In the more complex and realistic environment, the teacher intervenes when the student agent exhibits suboptimal behavior by assessing actions’ value to avert dangers. We also introduce an RL algorithm called Adaptive Clipping Proximal Policy Optimization (ACPPO), which combines samples from both teacher and student policies and employs dynamic clipping strategies based on sample importance. This approach improves sample efficiency while effectively alleviating data imbalance. Additionally, we employ the Kullback-Leibler divergence as a policy constraint, transforming it into an unconstrained problem with the Lagrangian method to accelerate the student’s learning. Finally, a gradual weaning strategy ensures that the student learns to explore independently over time, overcoming the teacher’s limitations and maximizing performance. Simulation experiments in highway lane-change scenarios show that the S2CD framework enhances learning efficiency, reduces training costs, and significantly improves safety compared to state-of-the-art algorithms. This framework also ensures effective knowledge transfer between teacher and student models, even with a suboptimal teacher, the student achieves superior performance, demonstrating the robustness and effectiveness of S2CD.
@@
Keywords Autonomous Vehicle  $\cdot$ Reinforcement Learning  $\cdot$ Knowledge Transfer  $\cdot$ Teacher-Student Framework  $\cdot$ Adaptive Clipping
@@ 1 Introduction
Due to the complexity and unpredictability of driving environments, approximately 94% of traffic accidents are related to suboptimal decisions made by human drivers (Alvaro et al., 2018). Suboptimal driving decisions not only compromise traffic safety but also reduce overall traffic efficiency (Lee et al., 2004). Hopefully, the rapid advancement of autonomous driving technology has attracted significant attention due to its potential to improve both traffic safety and efficiency. Autonomous driving systems generally consist of a perception layer, a decision-making layer, and a planning and control layer, in which the decision-making layer commonly referred to as the "brain" of the system (Chen et al., 2017). Ensuring the safety and accuracy of decision-making algorithms is therefore essential for the overall performance of autonomous driving systems (Li et al., 2017). Currently, decision-making methods for autonomous vehicles are generally classified into two categories: knowledge-driven and data-driven. Knowledge-driven methods, such as Hierarchical State Machines (HSM) (Patz et al., 2008), Expert Systems (ES) (Ferguson et al., 2008), and Finite State Machines (FSM) (Buehler et al., 2009), are known for their rigorous logic and high interpretability (Bae et al., 2020). However, these methods are limited by their strong reliance on existing knowledge, which restricts their ability to handle novel or unexpected scenarios (Bianco-Vega et al., 2005; Mozina et al., 2008). In contrast, data-driven methods such as End-to-End (E2E) Learning (Bojarski et al., 2016), Imitation Learning (IL) (Hawke et al., 2020), and Reinforcement Learning (RL) (Sallab et al., 2017; Zhang et al., 2022c) have gained prominence due to their robust self-learning capabilities and adaptability to dynamic environments. RL algorithms, in particular, optimize driving strategies through real-time interactions with the environment, making them highly effective in complex scenarios (Zhu and Zhao, 2021). While RL has demonstrated success in fields such as game artificial intelligence (AI) (Zhang et al., 2022a; Ye et al., 2020), financial transactions (Tsantekidis et al., 2023; Pendharkar and Cusatis, 2018), and energy management (Yu et al., 2019; Kuznetsova et al., 2013), its application in safety-critical areas like autonomous driving and robotic control faces significant challenges due to safety risks and training inefficiency (Liu et al., 2022; Ou et al., 2024).
@@
RL, much like human behavior, can autonomously explore environments and optimize strategies based on environmental feedback (Subramanian et al., 2022). However, traditional RL algorithms inevitably face risks during exploration, often encountering dangers before learning how to avoid them (Chow et al., 2018). The cost of allowing a vehicle to undertake dangerous actions in real-world scenarios is prohibitively high, making it challenging to apply traditional RL algorithms in autonomous driving. In human learning, particularly in hazardous situations, we do not rely solely on trial and error to acquire knowledge (Peng et al., 2022b). Instead, we seek guidance from a teacher to ensure safety and efficiency during the learning process. The teacher not only demonstrates correct actions when necessary but also directly corrects the student’s mistakes to avert dangerous situations. For instance, a driving instructor can take control of a vehicle when a novice driver makes an unsafe decision, thereby preventing accidents. Consequently, numerous researchers have incorporated the Teacher-Student Framework (TSF) (Zimmer et al., 2014; Liu et al., 2023) into RL, enabling student agents to safely and efficiently navigate complex scenarios through autonomous exploration and teacher guidance. In the TSF, a high-performance neural network or human expert acts as the teacher, intervening when the student’s actions meet specific criteria. When an intervention occurs, the teacher’s actions replace the student’s actions in the replay buffer as training data, which is later used to optimize the student’s model via RL algorithms. However, human-in-the-loop TSF requires continuous expert monitoring, which is labor-intensive, and the effectiveness of human teachers can be influenced by subjective factors like emotions and mental state. To overcome these issues, most TSF approaches rely on pre-training a high-performance teacher model using neural networks, allowing the student to learn correct behaviors from the teacher’s demonstrations. Yet, in fields like robotic control and autonomous driving, obtaining such a high-performance teacher in advance can be extremely expensive. Furthermore, in current TSF methods, if the teacher provides suboptimal demonstrations, the student may learn incorrect strategies, limiting its performance (Xue et al., 2023). Another limitation of current TSF methods is that the RL algorithm can only utilize the Markov Decision Process (MDP) corresponding to the action chosen by either the teacher or the student at each step, leading to significant data wastage since alternative actions are not leveraged.
@@
In summary, traditional RL algorithms involve high training costs but fail to guarantee the safety of agents during training. While standard TSF methods help student agents learn safely and efficiently, they come with significant costs in developing a high-performance teacher, and the student’s ultimate performance is constrained by the teacher’s abilities. Furthermore, these methods cannot optimize the policy by simultaneously utilizing samples generated by both the teacher and student models. To address these limitations, this paper introduces a novel TSF framework called Simple to Complex Collaborative Decision (S2CD), which enhances knowledge transfer between teacher and student. Unlike other TSFs, the S2CD framework accelerates training by first training the teacher model in a lightweight simulation environment with similar scenarios, then guiding the student in more complex simulations or real-world settings. We also theoretically demonstrated that teacher guidance improves both the performance and safety ceilings of the student during training. Additionally, we enhance the Proximal Policy Optimization (PPO) algorithm and propose Adaptive Clipping PPO (ACPPO), which utilizes data from both teacher and student models, significantly improving sample efficiency and alleviating data imbalance. The algorithm adaptively adjusts the clipping factor based on sample importance, boosting learning efficiency. During the policy optimization process, the neural network is updated by constraining the Kullback-Leibler (KL) divergence between the teacher’s and student’s policies, facilitating the rapid convergence of the student policy to the teacher’s. Finally, S2CD employs a weaning strategy that allows the student agent to independently explore the environment in later training stages, thus overcoming the teacher’s limitations. The main contributions of this paper are summarized as follows:
@@
  • This paper presents a novel TSF called S2CD, which accelerates the training of a teacher model in a lightweight simulation environment and uses it to guide the student agent in learning safely and efficiently in more complex simulated or real-world environments.
@@
  • This paper proposes the ACPPO algorithm, integrated within the S2CD framework, which simultaneously utilizes data from both teacher and student models while adaptively adjusting the clipping factor based on sample importance, thus enhancing sample efficiency and alleviating data imbalance.
@@
  • This paper introduces the use of the KL divergence between the teacher’s and student’s policies as a constraint during policy updates, transforming it into an unconstrained problem using the Lagrangian method to facilitate the rapid convergence of the student’s policy to the teacher’s.
@@
  • This paper investigates intervention and weaning strategies within the S2CD framework, where teacher interventions are triggered based on action values, and the frequency of interventions is gradually reduced to promote the student’s independent exploration, mitigating the limitations imposed by suboptimal teacher performance.
@@ 2.1 Knowledge Transfer
While RL has been widely applied across various fields and has demonstrated impressive results, most research has primarily focused on simulation experiments. However, in the context of autonomous driving, the ultimate goal is real-world application, where safety and efficiency during training process are critical factors in assessing algorithm performance. Training standard RL algorithms on real vehicles would not only risk substantial damage but also incur significant time costs. Knowledge transfer seeks to improve the learning efficiency of models by transferring knowledge from previously learned tasks to new tasks. By leveraging existing policies or experiences, the new task can accelerate the learning process, thereby reducing the need for extensive data collection or large sample sizes. This technique is especially valuable in environments where acquiring training data is costly or time-consuming. Consequently, effectively transferring knowledge from simulation-based training to agents operating in real-world or more complex simulation environments has become a significant focus of research.
@@
For example, Qiao et al. (2018) and Song et al. (2021) employed Curriculum Learning (CL) strategies to achieve notable performance in urban intersection and overtaking tasks, respectively. CL enhances model learning effectiveness by gradually increasing task difficulty, instead of requiring models to tackle complex tasks directly (Bengio et al., 2009). However, CL requires meticulously designed multi-level tasks. Furthermore, the progression from simpler to more complex tasks can increase training time and computational resource demands. Transfer Learning (TL) is another knowledge transfer technique that effectively reduces the training costs of the target task by applying knowledge acquired from the source task to the target task. For instance, Shu et al. (2021) combined RL with TL to transfer decision-making knowledge from one driving task to another in an intersection environment. Similarly, Akhauri et al. (2020) utilized simulated accident data for TL, enabling the model to generalize more rapidly and effectively to real-world scenarios. However, if the source and target tasks are not sufficiently related, the performance of TL may deteriorate, potentially resulting in "negative transfer" (Zhang et al., 2022b). Another approach to knowledge transfer is the use of Digital Twins (DT). By creating an accurate digital replica of a physical entity or system, we can map data from the physical world to the virtual world (Niaz et al., 2021), enabling real-time monitoring, simulation, analysis, and optimization of physical entities. Voogd et al. (2023) employed DT technology to combine virtual and real-world data on vehicle dynamics and traffic scenarios, effectively bridging the gap between simulation and reality. Wang et al. (2024) designed and implemented an E2E DT system for autonomous driving. This system captures real-world traffic information, processes it in the cloud to create a digital twin model, and then utilizes this model to provide route planning services for autonomous vehicles. Despite DT technology effectively bridging the gap between simulation training and real-world applications, accurate simulation processes remain indispensable. Designing simulation programs that accurately mirror or fully replicate real-world conditions is both challenging and costly (Yun and Park, 2021). Moreover, pre-training a high-performance model in a high-fidelity simulation environment is resource-intensive.
@@ 2.2 Safe RL
RL methods can enhance an agent’s ability to avoid dangers by learning from previous training failures and developing efficient behavioral strategies to adapt to complex environments. However, the application of these methods in safety-critical domains continues to present substantial safety challenges (Garcıa and Fernández, 2015). Ensuring policy safety throughout the training process and preventing agents from encountering dangerous situations are key to expanding the applicability of RL to more fields. Safe RL seeks to ensure that RL models avoid unsafe behaviors during both training and execution, particularly in high-risk or safety-critical domains such as autonomous driving. Safe RL techniques focus on balancing the exploration of novel strategies with the imperative to maintain safety, ensuring that the learned policies adhere to required safety standards throughout the learning process. Consequently, many researchers are focusing on incorporating constrained optimization, hard rule constraints, and risk assessment into RL algorithms to mitigate unsafe behaviors during both the learning and execution phases.
@@
For instance, methods like PPO-Lag (Stooke et al., 2020) and SAC-Lag (Ha et al., 2021) utilize the Lagrangian method, which simultaneously updates both the policy and the Lagrange multiplier through primal-dual optimization, thereby transforming the safety-constrained optimization problem into an unconstrained one. Peng et al. (2022a) proposed the Separated Proportional Integral Lagrangian (SPIL) algorithm, which mitigates oscillations and reduces conservatism in RL policies during car-following scenarios. Zhang et al. (2024) introduced cognitive uncertainty as a constraint in Lagrangian Safe RL algorithms, aiming to enhance both exploration and safety performance in autonomous vehicles. While the Lagrangian method is effective in handling safety constraints, it lacks a clear mechanism to prevent hazardous events (Li et al., 2022b). Thus, agents may still engage in unsafe exploration during training. Consequently, many researchers have applied rule-based constraints (Cao et al., 2022; Likmeta et al., 2020) or risk estimation methods (Mo et al., 2021; Li et al., 2022a) to Safe RL in autonomous driving, aiming to ensure a minimum level of policy safety performance. Although these methods can prevent a significant number of unsafe actions, their effectiveness heavily depends on the quality of the rule design or risk assessment models. To further reduce unsafe events, researchers have shifted their focus toward Safe RL approaches based on the TSF. Within the TSF, teacher policies can be classified into human expert strategies and well-trained neural network strategies. For example, Wu et al. (2023) and Huang et al. (2024) employed human experts to mentor AI agents. During the student agent’s interaction with the environment, the human expert intervenes in dangerous situations and demonstrates the correct actions, thereby ensuring driving safety throughout the student’s training process. While Safe RL methods involving human expert intervention significantly improve the safety of the exploration process, they incur high labor costs. Moreover, the teaching performance of human experts is highly influenced by subjective factors, leading to instability in the final model performance. To reduce labor costs, many researchers have employed well-trained neural network models as "teachers" to guide student agents during the learning process. For instance, Peng et al. (2022b) and Xue et al. (2023) employed RL algorithms to obtain a well-trained teacher model, when the conditions defined by a switching function, which determines whether to intervene, are met, the teacher takes over the student’s actions to demonstrate correct operations and avoid dangerous situations. Zhou et al. (2024) employed the SOAR cognitive architecture as a teacher to guide the student agent toward efficient and safe learning, while also offering a certain degree of interpretability.